• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name E.Chudakov

    Log entry time 23:59:28 on October 9,2006

    Entry number 186690

    keyword=Shift summary

    Beam came at about 17:00
    Up to 18:00 we checked what current the taget can hold with the cooling flow allowed and found that 80uA should be OK.

    18:00 - ARC started
    20:45 - Arun found a large beam spot. Accelerator experts are working on the problem.
    21:10 - Arun gave up with ARC measurement
    21:50 - beam restored (through the chicane), raster tuned

    RUNs
    1173 FPP doors open, 10cm He target, 20uA, T3=12k, 5M events
    1174 FPP 3+6", empty cell, 10uA, 2M events
    1175 Carbon, right sieve in, raster OFF, 80uA